A body moves along x axis such that it`s displacement varies with time as 7+3t+5t^2. The av velocity during 3rd sec

Grade:12th pass

3 Answers

Aditya Rana
6 years ago
38 m/sThe 3rd second displacement means the displacement in 4s minus the displacement in 3s. Putting the values into t, we get 99m in 4s and 61m in 3s. Subtracting them and dividing them by time which is 1s, we get 38.
Aditya Rana
6 years ago
I made a mistake, we have to find displacement in 3s and 2s ( because that is the third second, I answered for 4th second ), subtract them and divide by 1. So the answer is 28m/s
swarup kumar majee
6 years ago
displacement at time t=2 is x=7+6+20=33
at time t=3 displacement x=7+9+45=61
so average velocity in 3rd second is=61-33/1
=28m/s